OpenStudy (anonymous):

Derivative of (1/2)lnx?

OpenStudy (anonymous):

(1/2) is constant so it will come to left and the derivatic lnx

OpenStudy (campbell_st):

\[dy/dx = 1/2 \times 1/x = 1/2x\]